HUBBLE SPACE TELESCOPE DAILY REPORT #1773 PERIOD COVERED: 0000Z (UTC) 12/05/96 - 0000Z (UTC) 12/06/96 Daily Status Report as of 341/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed S/C/GHRS 6260 (Linear Extent and Ion Condition in = Lyman-Alpha Cloud: Cycle 5 Observations) The High Resolution Spectrograph was used to make spectrographic = observations between 1215 A0 and 1500 A0 of the QSO Q0107-025A in order = to study the linear extent and the ion condition in the Lyman-Alpha = cloud. The binary QSO system will be searched for any absorption systems = which may not be in common with the two objects, thus setting the limits = on the linear size of its clouds. A wavelength calibration image was = also taken. The observations were completed as planned and no anomalies = were reported. 1.2 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were reported.= 1.3 Completed FOS/BL 6913 (FOS Cycle 6: Location of Spectra and = Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make a = series of stability calibration measurements of the wavelength location = and the location of the spectra. All the observations were internals. = The observations were completed as planned, and no problems were = reported. 1.4 Completed GHRS/2 6599 (Boron in Halo Giants: A Solution to the D+3 = He Crisis in Big Bang Nucleus) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the boron abundances of giant halo star = HD63791. Comparison of the abundances among other stars of the same type = will allow the investigator to determine whether mass loss or mixing = causes the low Carbon-12/ Carbon-13 ratios seen in these type stars. Near = the beginning of this eight-and-a-half-hour proposal, at 340/122224Z, = during the primary and immediate re-acquisition, the scan steps limit was = exceeded. This parameter was back in limits at 340/122311Z. HSTAR 5827 = was written. Then, as described in HSTAR 5528, about an hour before the = end of the proposal, at 340/175746Z, the GHRS RIU B temperature slightly = exceeded the 29.0 deg. C limit. It remained out of limits until = approximately 340/1933Z. Otherwise, the observations executed as = scheduled, and no further problems were reported. 1.5 Completed FOS/BL 6781 (Soft X-Rays and the UV Spectra of a = Complete QSO Sample) The Faint Object Spectrograph (Blue Side) was used to obtain = ultraviolet spectra of the low-Z quasi-stellar object MARK-478. The = ionization continuum of this object has been studied using the ROSAT = satellite. The strengths and line profiles of strong and weak = ultraviolet emission and absorption lines and the ultraviolet continuum = were studied. The observations were completed as planned and no = anomalies were reported. 1.6 Completed FOS/RD 6577 (An Expanded Survey for Study of the = Lyman-Alpha Line and QSO Mg II Absorption) The Faint Object Spectrograph (Red Side) was used to make = observations of the QSO TB0933+733 in order to make a survey of the = Lyman-Alpha line and the 68 Mg II absorption line system. The = observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 8 Successful : 8 At 340/122224Z, during a primary and immediate re-acquisition, the = scan steps limit was exceeded. This parameter was back in limits at = 340/122311Z. HSTAR 5827 was written. 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: There were no reported operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH